Bayla works in a flower shop,where she produces q=8 floral arrangments per hour,h.She is paid 28 shekels an hour for the first nine hours she works and 35 shekels an hour for each additional hour.What is the firm's cost function,C(q)?What are its average cost (AC),average variable cost(AVC),and marginal cost(MC) functions?Draw the AC, AVC, and MC curves.


1
Expert's answer
2021-10-19T12:18:55-0400

Bayla earns 28 shekels per hour for the first 9 hours she works.


Therefore the firm's cost function will be "28\times9 =252" shekels.


She earns 35 shekels per hour for each additional hour she works. Assuming she works "x" hours after the initial 9 hours, her earnings will be "35\times x=35x" shekels.


  • Thus, the totals firm's cost function would be "TC=252+35x" shekels.


  • Average Cost "AC"="\frac{TC}{x}=\frac{252}{x}+35"


  • The average variable cost for the frim per hour will be "AVC=\frac{TVC}{x}=\frac{35x}{x}=35"


Now, the marginal cost for the firm could be obtained from the "TC" function. From the TC function,we can write;

"\delta TC=35\times\delta x"


Therefore,the marginal cost would be;

  • MC="\frac{\delta TC}{\delta x} =" 35 shekels.

AC, AVC and MC curves